# Quantum noise modeling through Reinforcement Learning

### v2

@@ IV.2 Quantum hardware
To test our algorithm on real quantum hardware, we used a single qubit in a superconducting transmon chip [70]. This 17 qubits chip has been produced by QuantWare  [^2] https://www.quantware.com and hosted at the Technology Innovation Institute of Abu Dhabi. The single qubit gate fidelity, obtained with RB, is $0.996$ with a readout fidelity of about $0.96$. To compute the DMs of the circuits, we used state tomography, running $4\times 10^{3}$ shots to compute each matrix. We attempted to mitigate measurement noise in advance to improve the fidelity of the DMs [71]. However, we observed that the RL agent performs better in learning the noise model without measurement error mitigation.
@@
To train the algorithm, we collected a dataset consisting of 60 random circuits of depth 10, employing $80\%$ of these circuits for the training set and the remaining $20\%$ for the test set. Figure 8 reports the average DM fidelity and TD during training for the first $2\times 10^{5}$ episodes. Convergence is reached after about $1.5\cdot 10^{5}$ episodes, reaching an average fidelity of $0.99$, with no evident signs of overfitting. This result is similar to the one obtained with simulations for single qubit circuits.
@@
Using a qubit with high gate fidelity makes the training process more challenging. The ground truth density matrices are affected by both gate errors, which are learned by the reinforcement learning (RL) algorithm, and by shot noise and measurement errors. In the high gate fidelity regime, measurement noise and shot noise can have an impact similar to that of gate noise. These errors in the density matrices make the reward signal less precise, thereby worsening the convergence of the training process. For our dataset, we have computed the average error on the trace distance introduced by measurement errors, shot noise and the additional gates needed to perform state tomography. The obtained value is $0.036$, which explains the large error bars observed during training, as shown in Figure 8. To mitigate this problem it would be useful to train the algorithm on longer circuits so that more gate errors can accumulate. In our case, to help the training process, it has been fundamental to set the maximum noise parameter value $P_{max}$ to a low value. As described in Section III.2, we have set $P_{max}$ to $0.008$, twice the decay parameter obtained with RB.
@@
Figure 9: Performance evaluation of different noise models on single qubit Clifford circuits of depths varying from 5 to 50. The performance is evaluated using average DM fidelity (left) and trace distance (right) with respect to the real noisy DMs obtained from execution on quantum hardware. The RL agent has been benchmarked against the RB method, the noiseless scenario, and the maximally mixed state (MMS). In the plot reporting the trace distance, we used a logarithmic scale to enhance distinguishability between RB and RL.
Noise Model 5 10 15 20 25 30 35 40 45 50
RL Avg 0.992 0.995 0.992 0.994 0.991 0.993 0.993 0.990 0.997 0.993
Std 0.009 0.005 0.005 0.003 0.005 0.003 0.002 0.008 0.002 0.005
RB Avg 0.992 0.994 0.992 0.992 0.987 0.990 0.992 0.981 0.992 0.990
Std 0.007 0.006 0.004 0.004 0.008 0.006 0.003 0.016 0.006 0.007
Table 2: Average fidelity and standard deviation in reconstructing noisy DMs of circuits with different depths obtained using the RL agent and RB noise model.
The performance benchmarking of the RL agent with respect to RB has been performed as described in Section IV.1. For this test, we used circuits of length spanning from 5 to 50, 10 circuits for each length. As the noise level of the qubit is quite low, it is necessary to use circuits with higher lengths to extract the decay parameter for the RB method. Figure 9 reports the performance, in reconstructing the noisy DMs, of the trained RL agent compared with RB. The comparison also includes circuits without noise and the maximally mixed state. To better observe the performance differences between the RL and RB noise models, we used a logarithmic scale in the plot reporting the TD. On average, the RL agent outperforms the RB method for all circuit lengths when using the trace distance as the metric. The high standard deviation obtained in the plot is mainly due to the uncertainty in the evaluation of the DMs with state tomography. Regarding the fidelity, we present the results obtained with RB and RL in Table 2, as the two series are not easily distinguishable from the plot. For this metric as well, the performance of the RL agent is better than or equal to RB for all circuit lengths.
@@
The results obtained in this section differ from those obtained in Section IV.1, where the RL agent clearly outperformed the RB method. However, it is important to highlight that the RL agent requires fewer hardware resources than the RB method. In the training of the RL agent, a dataset of 60 circuits with a depth of 10 has been sufficient. The dataset generation process on the employed hardware took just a few minutes, at the cost of some classical resources needed for training the algorithm. Conversely, to estimate the RB parameter with good precision, it is necessary to run many circuits for each depth.
V Applications
@@
This section we assess our model’s performance on Quantum Fourier Transform (QFT) and Grover’s algorithm circuits under simulated noise. These tests provide valuable insights into the generalization capabilities of the model and serve as a stress test and benchmark for overall performance.
@@
$\operatorname{QFT}\ket{x}=\frac{1}{\sqrt{2^{n}}}\sum_{k=0}^{2^{n}-1}e^{2\pi ixk/2^{n}}\ket{k}\,.$ (6)
Being a unitary transformation, the QFT can be implemented as a quantum circuit. For a three-qubit system ($n=3$), the QFT circuit is reported in Figure 10.
{quantikz}
@@
Figure 10: QFT circuit for three qubits. In this circuit, $R_{k}=U_{1}(2\pi/2^{k})$ and $H$ denotes the Hadamard gate.
Grover’s algorithm, another cornerstone of quantum computing, is renowned for its ability to search unsorted databases with quadratic speedup compared to classical algorithms [3]. The algorithm operates on a superposition of quantum states, and its goal is to find a specific state |$\ket{w}$ that satisfies a certain condition defined by an oracle function. The key component of Grover’s algorithm is the Grover iterate, a unitary transformation that contain the information of the oracle. The Grover iterate is typically repeated $\mathcal{O}{\sqrt{N}}$ times to maximize the probability of measuring |$\ket{w}$, where $N$ is the dimension of the system. We considered a two-qubit system and the target state |$\ket{11}$. This configuration requires only one Grover iteration to find the target state. We utilized an ancillary qubit to construct the oracle, leading to the final circuit reported in Figure 11.
{quantikz}
[…]
Figure 11: Grover’s search algorithm circuit, the target state is |$\ket{11}$ and an ancillary qubit is required.
We transpiled the circuits to utilize the native gates $R_{z}$, $R_{x}$, and $CZ$. The gate number after transpilation and other significant circuit parameters are detailed in Table 3. It should be noted that the increased depth of the circuits, the lower fraction of two-qubit gates, and the structure, compared to the circuits in the training set, make this generalization test particularly challenging.
Circuit Total gates $CZ$ gates Moments
@@
Table 3: Number of gates and circuit moments of the transpiled circuits for QFT and Grovers’s algorithm.
We have tested the algorithms using both the high and low noise models for thee-qubit circuits detailed in Section IV.1 This approach served a dual purpose. Primarily, we aimed to test the algorithm’s generalization capabilities with noise models exhibiting lower error rates. Secondly, the noise model from Section IV.1 completely masked the final result for Grover’s algorithm. By using a noise model with lower error rates, we ensured that a peak at the target state in the final result remained discernible. We evaluated the performance of our RL agent against the RB noise model and the limit case where no noise is added to algorithm’s circuits. The fidelity between the reconstructed and original noisy DMs for the different noise models is detailed in Table 4. In all instances, the RL agent obtained the best performance.
@@
Table 4: Fidelity between the density matrix reconstructed with a noise model (RL agent, RB, the limit case where no noise channels are added and the MMS) and the ground truth noisy one. The result is reported for QFT and Grover’s algorithm circuits for both a high and low noise models.
Figure 12: Computational basis states probabilities for the QFT circuit with a high error noise model (top) and Grover’s algorithm circuit with a low error noise model (bottom). The histograms show a comparison between probabilities obtained with the ground truth noise model, the RL agent noise model and the RB noise model.
Figure 13: Heatmap of the absolute error between the ground truth noisy DM and the DM obtained with the RL agent (left) and RB model (right) for the three-qubit QFT circuit.
The final state probabilities for both the QFT and Grover’s algorithm circuits, measured in the computational basis, are shown in Figure 12. For the QFT circuit we report the result obtained with a high error noise model while for the Grover’s circuit we report the result obtained with a low error noise model that doesn’t destruct the expected result. The histograms compare the probabilities derived from the original noisy circuit, the reconstructed circuit using the RL agent, and the RB noise model. Given the depth of the circuits used, the RB noise model tends to average the output. In contrast, the RL agent, with a few exceptions, aligns more closely with the probabilities obtained from the original noise model. This alignment is particularly noticeable in the QFT simulation counts for the state |$\ket{000}$. This state has the higher probability due to the noise model’s amplitude damping channels, a feature that is successfully replicated by the RL agent but that is is not possible to replicate with the RB model.
For a more detailed analysis, Figure 13 shows the absolute error between the DMs generated by the RL algorithm and the RB model compared to the ground truth noisy DM. The errors in the RL model are well distributed across the DM, while the RB model’s errors tend to be higher along the diagonal. This effect was observed in many circuits, even during preliminary tests with the RL algorithm. As the depth of the circuit increases, the density matrix of noisy circuits tends to approach the MMS, resulting in most of the information being contained in the diagonal. This is one of the main reasons we chose the trace distance as the metric for the reward of the RL algorithm.
The results obtained in this section underscore the generalization capabilities of the proposed RL approach for noise modeling. The RL agent adapts to circuits with structures distinct from the random ones used in the training set and of significantly different depths. This allows the algorithm to be used for circuit simulations for interesting use cases like quantum algorithms and quantum machine learning.
@@ VI Conclusions
This work presents a reinforcement learning algorithm for replicating specific noise models in single and multiple qubit quantum circuits. This approach reduces heuristic assumptions about the noise model, enhancing generalization properties. Tests on simulated and quantum hardware circuits have demonstrated the model’s ability to learn complex noise patterns and generalize to unseen circuits. For a comprehensive evaluation and to show a possible use case, we tested the model on QFT and Grover’s algorithm circuits. In all occasions, the RL model consistently outperformed a common noise characterization method, randomized benchmarking both in the ability of reconstructing the density matrices and in the amount of quantum hardware resources needed.
Possible future applications of the algorithm include not only reproducing the noise pattern of a specific hardware device. By learning the error patterns of qubits for specific gate types, the model could optimize the transpilation process [72], thereby enhancing quantum algorithms fidelity. Furthermore, using the knowledge of the noise for its mitigation could be an interesting approach.
The current model’s limitation is its scalability to circuits with many qubits. There are two main challenges to overcome. The first problem is that scaling the model would necessitate a significant increase in the number of actions, A lager action space would complicate and slow down the training process. The second problem regards obtaining the ground truth density matrices via quantum state tomography as it requires exponentially more measurements in the number of qubits. We are considering potential solutions to these challenges. One approach, to solve the second issue, could involve training the model with probability distributions derived from measurements, rather than density matrices. It is also possible to employ machine learning techniques to reduce the amount of measurements needed for quantum state tomography [73, 74]. To address the first issue, we could partition large circuits into smaller ones, facilitating parallel training of multiple smaller models. Furthermore, it could be beneficial to explore the use of graph neural networks in place of convolutional neural networks to encode qubit connectivity information into the model.
@@
Author contribution: S. Bordoni, A. Papaluca, P. Buttarini, A. Sopena: Methodology, Software, Writing.
S. Giagu, S. Carrazza: Project Administration, Review, Funding.
Funding: This work is partially supported by the Technology Innovation Institute (TII), Abu Dhabi, UAE.
S. B. and S. G. are partially supported by ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union - NextGenerationEU.
A. S. is supported through the Spanish Ministry of Science and Innovation grant SEV-2016-0597-19-4, the Spanish MINECO grant PID2021- 127726NB-I00, the Centro de Excelencia Severo Ochoa Program SEV-2016-0597 and the CSIC Research Platform on Quantum Technologies PTI-001.
@@
Conflict of interest: The authors declare no conflict of interest.
Open access: This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author(s) and the source, provide a link to the Creative Commons license, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons license, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons license and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder. To view a copy of this license, visit https://creativecommons.org/licenses/by/4.0/.

### v3

@@ IV.2 Quantum hardware
Figure 7: Average density matrix fidelity (left) and trace distance (right) throughout the training process of the RL agent on single-qubit circuits executed on a superconductive quantum chip. The metrics have been evaluated on a dataset of 60 circuits using 80% for the training set and 20% for the test set; shaded regions represent the standard deviation. Convergence was achieved after approximately $2\times 10^{5}$ episodes, reaching an average fidelity of $0.99$ on the test set. No evident signs of overfitting were observed.
@@
To test our algorithm on real quantum hardware, we used a single-qubit in a superconducting transmon chip [21]. This 17 qubits chip has been produced by QuantWare  [^4] https://www.quantware.com and hosted at the Technology Innovation Institute of Abu Dhabi. The single-qubit gate fidelity, obtained with RB, is $0.996$ with a readout fidelity of about $0.96$. To compute the DMs of the circuits, we used state tomography, running $4\times 10^{3}$ shots to compute each matrix. We attempted to mitigate measurement noise in advance to improve the fidelity of the DMs [54]. However, we observed that the RL agent performs better in learning the noise model without measurement error mitigation.
@@
To train the algorithm, we collected a dataset consisting of 60 random circuits of depth 10, employing $80\%$ of these circuits for the training set and the remaining $20\%$ for the test set. Figure 7 reports the average DM fidelity and TD during training for the first $2\cdot 10^{5}$ episodes, after this episode no performance improvements have been observed. At convergence, the model reached an average fidelity of $0.99$, with no evident signs of overfitting. This result is similar to the one obtained with simulations for single-qubit circuits.
@@
Using a qubit with high gate fidelity makes the training process more challenging. The ground truth density matrices are affected by both gate errors, which are learned by the RL algorithm, and by shot noise and measurement errors. In the high gate fidelity regime, measurement noise and shot noise can have an impact similar to that of gate noise. These errors in the density matrices make the reward signal less precise, thereby worsening the convergence of the training process. For our dataset, we have computed the average error on the trace distance introduced by measurement errors, shot noise, and the additional gates needed to perform state tomography (see Appendix B). The obtained value is $0.036$, which explains the large error bars observed during training, as shown in Figure 7. To mitigate this problem it would be useful to train the algorithm on longer circuits so that more gate errors can accumulate. In our case, to help the training process, it has been fundamental to set the maximum noise parameter value $P_{max}$ to a low value. As described in Section III.2, we have set $P_{max}$ to $0.008$, twice the decay parameter obtained with RB.
@@
Figure 8: Performance evaluation of different noise models on single-qubit Clifford circuits with depths ranging from 5 to 50. The performance is assessed using average DM fidelity (left) and trace distance (right), with respect to the real noisy DMs obtained from execution on quantum hardware. The RL agent is benchmarked against the RB method, the noiseless scenario, and the maximally mixed state (MMS). In the trace distance plot, a logarithmic scale is used to enhance the distinguishability between RB and RL. The lower level of noise, compared to the simulation, can be inferred from the fidelity with the MMS, which does not reach the same values as in the simulation. A comparison between the RL and RB models shows that the former outperforms the latter on average; however, due to large errors, their performances are comparable within one standard deviation.
The performance benchmarking of the RL agent with respect to RB has been performed as described in Section IV.1. For this test, we used circuits of length spanning from 5 to 50, 10 circuits for each length. As the noise level of the qubit is quite low, it is necessary to use circuits with higher lengths to extract the decay parameter for the RB method.
Figure 8 summarizes the performance of the trained RL agent in reconstructing noisy density matrices, compared with the RB method. The comparison also includes results for circuits without noise and for the maximally mixed state. To highlight differences between the RL and RB noise models, we evaluated the TD using a logarithmic scale. On average, the RL agent outperforms the RB method across all circuit lengths, although their performances are statistically comparable within one standard deviation. The relatively large standard deviation observed is primarily due to the limited number of circuits in the test set and the presence of measurement errors, which introduce additional uncertainty in the ground truth DMs.
@@
The results obtained in this section differ from those obtained in Section IV.1, where the RL agent clearly outperformed the RB method. However, it is important to highlight that the RL agent requires fewer hardware resources than the RB method. In the training of the RL agent, a dataset of $60$ circuits with a depth of $10$ has been sufficient (for a total number of $600$ circuit moments). The dataset generation process on the employed hardware took just a few minutes, at the cost of some classical resources needed for training the algorithm. Conversely, to estimate the RB parameter with comparable precision, it is necessary to run a larger number of circuits. In our case, we needed to run a total of $110$ circuits with an average depth of $27.5$, for a total number of $3025$ circuit moments. Performing the RB protocol required approximately five times more quantum hardware resources than generating the training set for the RL agent.The overhead given by the RL training process is negligible because it is performed on classical hardware and does not require extensive resources. The NN used for the algorithm has a total number of parameters of about $10^{4}$, which does not require a large amount of memory to be stored. The training process can be performed on a standard laptop in minutes without even requiring a GPU. Scaling the algorithm to circuits with a higher number of qubits may be a challenge as a bigger action space and observation space would require a larger NN to be trained. Addressing this challenge is a potential direction for future work. One possible solution is to divide the agent into multiple sub-agents, each responsible for a subset of the qubits.
It is interesting to note that calibration parameters, as well as other external factors that may affect the noise levels of a quantum chip, tend to vary over time. This implies that the RL agent should be retrained regularly on a new dataset to accurately capture the current noise conditions. This issue affects all noise modeling techniques. It is difficult to define a fixed time window during which the RL model, or any noise model, remains valid. This validity period depends on various factors, including the characteristics of the chip, refrigerator, electronics, and even the external conditions of the laboratory. Under stable conditions, the model can be considered valid until a new calibration routine is performed, which typically occurs on a daily basis.
V Test on quantum algorithms
@@
In this section we assess our model’s performance on the Quantum Fourier Transform (QFT) and Grover’s algorithm circuits under simulated noise. These tests provide valuable insights into the generalization capabilities of the model and serve as a stress test and benchmark for overall performance.
@@
Being a unitary transformation, the QFT can be implemented as a quantum circuit. For a three-qubit system ($n=3$), the QFT circuit is shown in Figure 9.
             $H$ $H$ $R_{2}$ $R_{2}$ $R_{3}$ $R_{3}$ $H$ $H$ $R_{2}$ $R_{2}$ $H$ $H$
Figure 9: QFT circuit for three qubits. In this circuit, $R_{k}=R_{z}(2\pi/2^{k})$ and $H$ denotes the Hadamard gate. The last SWAP gate can be omitted as this reordering can be handled classically. This circuit implementation has to be transpiled into native gates before being passed to the RL agent.
@@
We transpiled the circuits to utilize the native gates $R_{z}$, $R_{x}$, and $CZ$. The gate number after transpilation and other significant circuit parameters are detailed in Table 3. A random non-Clifford circuit of the same type used for the training set is also included as a benchmark. It should be noted that the circuits implementing the QFT and Grover algorithms are substantially different from those used in the training set. The main differences include increased circuit depth, a lower fraction of two-qubit gates, and structural variations. Some gate patterns present in these circuits were likely never encountered by the agent during training, as they would be unlikely to arise through random circuit construction. All these factors contribute to making the generalization test performed in this section particularly challenging.
Circuit Total gates CZCZ gates Depth
@@
Random 23 12 15
Table 3: Number of gates and circuit depth of the transpiled circuits for QFT and Grovers’s algorithm. A random circuit from the training set is also included as a benchmark to highlight the differences including depth and a higher fraction of two-qubit gates.
We have tested the algorithms using both the high and low noise models for three-qubit circuits detailed in Section IV.1. This approach served a dual purpose. Primarily, we aimed to test the algorithm’s generalization capabilities with noise models exhibiting lower error rates. Secondly, the noise model from Section IV.1 completely masked the final result for Grover’s algorithm. By using a noise model with lower error rates, we ensured that a peak at the target state in the final result remained discernible. We evaluated the performance of our RL agent against the RB noise model and the limit case where no noise is added to algorithm’s circuits. A comparison with the maximally mixed state is also given to give an estimate of the total level of noise. The fidelity between the reconstructed and original noisy DMs for the different noise models is detailed in Table 4. We also report the result obtained with the random circuit as a benchmark. In all instances, the RL agent obtained the best performance.
@@
Random High 0.99 0.93 0.61 0.62
Low 0.99 0.98 0.77 0.51
Table 4: Fidelity between the density matrix reconstructed with a noise model (RL agent, RB, the limit case where no noise channels are added and the MMS) and the ground truth noisy one. The result is reported for QFT and Grover’s algorithm circuits for both a high and low noise models. The result obtained with a non-Clifford random circuit is also reported for comparison. The RL agent outperforms, or is comparable, with respect to the RB model in all the studied cases.
Figure 11: Computational basis states probabilities for the QFT circuit with a high error noise model (top), Grover’s algorithm circuit with a low error noise model (mid) and a random non-Clifford circuit with a high error noise model (bottom). The histograms show a comparison between probabilities obtained with the ground truth noise model, the RL agent noise model and the RB noise model. It is possible to observe that while the RB model tends to flatten the distribution, the RL agent is able to reconstruct the peaks with more fidelity.
Figure 12: Heatmap of the absolute error between the ground truth noisy DM and the DM obtained with the RL agent (left) and RB model (right) for the three-qubit QFT circuit. It is possible to observe that the RL agent is able to better reconstruct the noisy DM not only in the diagonal part. This behavior has been observed in all the circuits tested in this section.
[…]
For a more detailed analysis, Figure 12 shows the absolute error between the DMs generated by the RL algorithm and the RB model compared to the ground truth noisy DM. The errors in the RL model are well distributed across the DM, while the RB model’s errors tend to be higher along the diagonal. This effect was observed in many circuits, even during preliminary tests with the RL algorithm. As the depth of the circuit increases, the density matrix of noisy circuits tends to approach the MMS, resulting in most of the information being contained in the diagonal.
The results obtained in this section underscore the generalization capabilities of the proposed RL approach for noise modeling. The RL agent adapts to circuits with structures distinct from the random ones used in the training set and of significantly different depths. This allows the algorithm to be used for circuit simulations for interesting use cases like quantum algorithms and quantum machine learning when the exact noise model of the hardware chip is unknown.
@@ VI Conclusions
This work presents a novel reinforcement learning algorithm for accurately replicating complex noise models in single- and multi-qubit quantum circuits. By learning noise patterns directly from experimental data, this approach significantly reduces heuristic assumptions about the underlying noise model, thereby enhancing generalization properties. Tests on simulated and quantum hardware circuits have demonstrated the model’s ability to learn complex noise patterns and generalize to unseen circuits. For a comprehensive evaluation and to show a possible use case, we tested the model on QFT and Grover’s algorithm circuits. In all cases, the RL model consistently outperformed a common noise characterization method, randomized benchmarking both in the ability of reconstructing the density matrices and in the amount of quantum hardware resources needed.
Possible future applications of the algorithm include not only reproducing the noise pattern of a specific hardware device. By learning the error patterns of qubits for specific gate types, the model could optimize the transpilation process [91], thereby enhancing quantum algorithms fidelity. Furthermore, using the knowledge of the noise for its mitigation could be an interesting approach. In this direction, it would be useful to employ explainable machine learning techniques to gain insights into the noise model learned by the RL agent.
The primary limitation of the current model is its scalability to quantum circuits with many qubits. This challenge manifests through two fundamental bottlenecks that must be addressed for practical deployment. The first challenge concerns the growth of the action space. Scaling the model to larger qubit systems requires an increase in the number of possible actions, resulting in a larger expanded action space that complicates the training process. The second challenge relates to the acquisition of ground truth density matrices via quantum state tomography, which requires an exponentially increasing number of measurements with respect to the number of qubits, making the approach intractable for large systems. For the action space scalability issue, we propose partitioning large circuits into smaller, manageable subsystems, enabling parallel training of multiple specialized smaller models. Furthermore, replacing convolutional neural networks with graph neural networks could prove beneficial, as they can naturally encode qubit connectivity information and circuit topology into the model architecture. To address the tomography bottleneck, one approach involves training the model directly on probability distributions derived from computational basis measurements, rather than full density matrices. Additionally, advanced machine learning techniques can be employed to reduce the measurement overhead required for quantum state tomography [43, 78].
@@
Author contribution: S. Bordoni, A. Papaluca, P. Buttarini, A. Sopena: Methodology, Software, Writing.S. Giagu, S. Carrazza: Project Administration, Review, Funding.
Funding: This work is partially supported by the Technology Innovation Institute (TII), Abu Dhabi, UAE.S. B. and S. G. are partially supported by ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union - NextGenerationEU.A. S. was supported through the Spanish Ministry of Science and Innovation grant SEV-2016-0597-19-4, the Spanish MINECO grant PID2021- 127726NB-I00, the Centro de Excelencia Severo Ochoa Program SEV-2016-0597 and the CSIC Research Platform on Quantum Technologies PTI-001.
@@
Open access: This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author(s) and the source, provide a link to the Creative Commons license, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons license, unless indicated otherwise in a credit line to the material.
Appendix A Randomized benchmarking as a noise model
The decay parameter $f$, extracted through randomized benchmarking (RB), can be used to construct a simple noise model for the quantum device. Consider one of the RB circuits consisting of $m$ gates. Let $\rho$ denote the density matrix of the initial pure state, and assume that each gate introduces a depolarizing channel characterized by a parameter $\lambda\in\quantity(0,1)$. The action of the overall noise channel $\mathcal{E}$ on the state after $m$ gates is
$\mathcal{E}(\rho)=(1-\lambda)^{m}\rho+\quantity(1-(1-\lambda)^{m})\frac{I}{d}$ (A1)
where $d$ is the dimension of the system. The fidelity between the noisy state $\mathcal{E}(\rho)$ and the ideal state $\rho$ is given by
$F(\mathcal{E}(\rho),\rho)=(1-\lambda)^{m}+\frac{1-(1-\lambda)^{m}}{d}\,.$ (A2)
[…]
Appendix B Uncertainty estimation for the trace distance
We estimate the error in the trace distance $T(\rho,\sigma)$ between the noisy density matrix $\rho$, obtained via state tomography, and the model-generated density matrix $\sigma$, by propagating the uncertainties of each element of $\rho$ as
$\Delta T(\rho,\sigma)=\sum_{i,j}\left|\frac{\partial T}{\partial\rho_{ij}}\right|\Delta\rho_{ij}\,$, (A3)
where $\Delta\rho_{ij}$ is the uncertainty in the matrix element $\rho_{ij}$. The partial derivatives are approximated using finite differences.
We consider three main sources of error: (i) statistical fluctuations due to the finite number of shots, (ii) measurement errors, and (iii) errors introduced by additional gates required for basis changes in state tomography. Measurement errors are modeled as single-qubit bit-flip channels with probability $1-f_{r}$, where $f_{r}$ is the readout fidelity. Errors from basis-change gates are modeled as depolarizing noise with parameter 1f1-f, as described in the main text.
A single-qubit density matrix can be written as
$\rho=\frac{1}{2}(I+\vec{r}\cdot\vec{\sigma})\,$, (A4)
where $\vec{r}=(r_{x},r_{y},r_{z})$ is the Bloch vector and $\vec{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are the Pauli matrices. The components of the Bloch vector are given by
$r_{j}=\text{Tr}(\sigma_{j}\rho)\,,\quad\text{for }j=x,y,z\,.$ (A5)
Assuming $\rho$ is a noisy density matrix, the uncertainty on the Bloch vector components can be expressed as
[…]
where $N_{s}$ is the number of shots, $a_{j}$ is the expectation value of $\sigma_{j}$ for the noiseless density matrix, and $\tilde{\rho}_{j}$ is the density matrix in basis $j$ including readout noise. The uncertainty on the density matrix elements is then given by
$\Delta\rho_{ij}=\frac{1}{2}\vec{\Delta r}\cdot\vec{\sigma}\,$, (A7)
where $\Delta\vec{r}=(\Delta r_{x},\Delta r_{y},\Delta r_{z})$ is the vector of uncertainties on the Bloch vector components.